find y=2x^6 - 10x^3 + 2x, find d^5 y/ dx^5

d^5 y/ dx^5=
Answer #1

y = 226_1ox? + de d (23€ _ 10x3 + 2x) z 12x5T - 304272 oder einen (12x®- 3042 +23 = 6024 - 60% da = a hal boxt-boxs 15 = 240